# Bitcoin Under Volatile Block Rewards: How Mempool Statistics Can Influence Bitcoin Mining

### v2

@@ D.3. Strategy $\pi_{2}^{\text{Non-predictable}}$
Let $P_{0}$, $P_{1}$, $P_{2}$ denote the stationary probability of being in states $S_{0}$, $S_{1}$, and $S_{2}$, respectively. We have:
(15) $\begin{split}&P_{0}=1-\left(1-\alpha^{2}\right)p,\quad P_{1}=\left(1-\alpha\right)p,\quad\\ &P_{2}=\alpha\left(1-\alpha\right)p.\end{split}$
@@
The normalized average number of canonical honest blocks, canonical adversarial blocks, can be obtained as follows:
@@
(16) $\begin{split}N_{\mathcal{H}}=&P_{0}\left(1-\alpha\right)\left(1-p\right)+P_{1}\left(1-\alpha\right)p+2P_{1}\left(1-\alpha\right)\left(1-p\right)+\\ &P_{2}\left(1-\alpha\right)p+2P_{2}\left(1-\alpha\right)\left(1-p\right),\\ N_{\mathcal{A}}=&P_{0}\alpha+2P_{2}\alpha,\\ R_{\mathcal{A}}=&P_{0}\alpha\left(1+Fp\right)+P_{2}\alpha\left(2+F(1+p)\right)\enspace.\end{split}$
@@ Appendix E Limitations of the Simplified Environment
Non-linearity of the time-fee equation: In Fig. 4 presented in Section 4.1, a linear regression model is used to represent the time-fee equation. However, the block transaction fee does not necessarily increase linearly with the block generation time. This occurs when the rate of fee increase is not constant and decreases as the block generation time increases. In such cases, using alternative regression techniques rather than a linear model can provide a time-fee equation that better fits the behavior observed in the Bitcoin mempool. In Fig. 13(a), we compare linear regression with a curve regression model given by $f(t)=0.6414\;\texttt{ln}(t+6.5209)-0.8419$ to derive the time-fee equation for Bitcoin blocks mined on May 3, 2023. The coefficient of determination (R-squared) for the curve regression is $0.6243$, while for the linear regression it is $0.5180$. This indicates that the curve regression provides a better fit compared to the linear regression, suggesting that the fee increase rate diminishes with longer block generation times.
Dependence of fee rewards on the generation times of preceding blocks: Another technique that can lead to a more realistic mempool implementation is using multivariate regression instead of univariate regression. In univariate regression, the block transaction fee is estimated based solely on a single predictor, namely the block generation time. In contrast, multivariate regression allows us to account for additional predictors, such as the generation time of the parent block, in determining the block transaction fee. In Fig. 13(b), we depict the regression plane for the block transaction fee on March 18, 2024, based on two predictors: the block generation time and the parent block generation time. The resulting time-fee equation is $f(t_{1},t_{2})=0.1697+0.0079*t_{1}+0.0046*t_{2}$, where $t_{1}$ and $t_{2}$ represent the block generation time and the parent block generation time, respectively. Note that if the parent block generation time is longer, more fee-valuable transactions will likely remain in the mempool after the parent block is mined, allowing the next block to collect these transactions. As the time-fee equation shows, the fee increase rate with respect to the block generation time is higher than the fee increase rate with respect to the parent block generation time. This implies that the block generation time is a more effective predictor of the transaction fee compared to the parent block generation time.
@@
Figure 13. Time-fee equation
@@ Appendix F A3C Implementation
F.1. A3C Actions and State Representation
$\texttt{adopt}_{i}$: The action $\texttt{adopt}_{i}$, $i\in\{0,1,\ldots,k_{1}\}$ represents that the adversary abandons its adversarial fork and starts mining a new fork on top the longest public chain whose latest $i$ blocks are pruned. In the fixed block reward model, considering only the action $\texttt{adopt}_{0}$—namely adopting the longest chain—is enough as $\texttt{adopt}_{i}$ for $i>0$ cannot lead to a more profitable strategy. However, in the volatile block reward model, a deep block in the honest fork may have an extraordinary amount of transaction fees, incentivizing the adversary to orphan the block and steal its transaction fees. A block with an extraordinary fee can result from including a whale transaction as considered in (Bar-Zur et al., 2023) or from having a long generation time. Action $\texttt{adopt}_{i}$ is feasible if $i\leq l_{\mathcal{H}}$.
$\texttt{undercut}_{\texttt{block}}$: This action represents that the adversary tries to undercut the tip block of the longest public chain denoted by $B^{\text{tip}}_{\mathcal{H}}$. To perform block undercutting, the adversary: i) abandons its fork, ii) adopts $l_{\mathcal{H}}-1$ blocks in the honest fork, namely the entire honest fork except for the last block $B^{\text{tip}}_{\mathcal{H}}$, and iii) adjusts the total transaction fee in its potential next block to be less than the total transaction fee included in $B^{\text{tip}}_{\mathcal{H}}$. If, after taking the action $\texttt{undercut}_{\texttt{block}}$, the next block is mined by the adversary, the adversary’s subsequent action is match. Choosing action match will result in a race between two forks of the same height, differing only in their last block. In this scenario, the altruistic miners will continue mining on top of the honest fork since they received it earlier. However, the petty-compliant miners will switch to mining on top of the adversarial fork as it contains a lower amount of transaction fees. Our implementation considers the $\epsilon$-petty-compliant miners with an adjustable $\epsilon$, where the petty-compliant miners deviate from the honest strategy if doing so allows them to earn at least $\epsilon$ BTC more. If the amount of remaining transaction fee in the transaction pool is relatively low and the tip block $B^{\text{tip}}_{\mathcal{H}}$ includes a relatively large amount of transaction fee, undercutting block $B^{\text{tip}}_{\mathcal{H}}$ may be a logical action for the adversary as it enables the adversary to fill its block with transactions with higher fees. This action is always feasible.
In the state representation above, $l_{\mathcal{A}}$ ($l_{\mathcal{H}}$) denotes the adversarial (honest) fork length. Latest is a boolean variable that represents whether the latest mined block is honest or adversarial. If the last block is honest and $l_{\mathcal{A}}\geq l_{\mathcal{H}}$, the adversary can take action match. Match is a boolean variable that represents whether the same-height fork race resulting from the action match is resolved or not. If after taking action match, the next block is honest, the fork race is resolved. However, if the next block is adversarial, the fork race persists. undercut is a boolean variable that represents whether the next block mined after taking actions $\texttt{undercut}_{\texttt{block}}$ or $\texttt{undercut}_{\texttt{fork}}$ is adversarial or not. If the adversary manages to mine the next block, it should then take action match to publish the undercutting block or fork to the rational miners. $BR_{\mathcal{A}}^{i}$ for $i\in\{1,2,\ldots,\texttt{maxForkLen}\}$ denotes the block reward of the $i^{\text{th}}$ block in the adversarial fork. This block reward includes both the protocol reward and the transaction fees collected from the included transactions. Note that maxForkLen is a variable that determines the maximum length of a fork in our implementation. $\texttt{Pool}_{\mathcal{A}}^{i}$ ($\texttt{Pool}_{\mathcal{H}}^{i}$) for $i\in\{1,2,\ldots,\texttt{maxForkLen}\}$ denotes the mempool statistics w.r.t. the chain whose tip block is the $i^{\texttt{th}}$ adversarial (honest) block. Once a new block is mined, it includes some of the transactions available in the mempool, resulting in the removal of those transactions from the mempool and consequently a modification in the mempool statistics. A state must store the mempool statistics w.r.t. the $i^{\texttt{th}}$ adversarial (honest) block as miners may need to abandon their current forks and continue mining on top of the $i^{\texttt{th}}$ adversarial (honest) block as a result of a specific action. Mempool statistics is an $N^{\text{range}}$-element array that contains the weight of transactions available in each of the satoshi-per-byte ranges $\{\texttt{fee}_{0},\texttt{fee}_{1},\ldots,\texttt{fee}_{N^{\text{range}}-1}\}$. The weight of transactions corresponding to each satoshi-per-byte range $\texttt{fee}_{j}$ increases over time according to the specific weight-time equation defined for the range $\texttt{fee}_{j}$. As mentioned earlier, once a new block is mined, a total weight of 1 virtual megabyte is deducted from the transactions in the highest satoshi-per-byte ranges. However, if the adversary takes one of the actions $\texttt{undercut}_{\texttt{block}}$ or $\texttt{undercut}_{\texttt{fork}}$, the block may be filled with less than 1 virtual megabyte of transactions to leave some valuable transactions in the mempool for petty-compliant miners. $\texttt{Pool}_{\mathcal{A}}^{i}$ ($\texttt{Pool}_{\mathcal{H}}^{i}$) is an array containing $N^{\text{range}}$ elements, where the $j^{\text{th}}$ element represents the weight of total transactions of the mempool that pay $\texttt{fee}_{j-1}$ satoshis per virtual byte. Finally, $\texttt{Pool}_{\mathcal{C}}$ denotes the mempool statistics w.r.t. the canonical chain.
F.2. Objective Function
@@
F.2.1. Before the difficulty adjustment
As discussed in Section 2.3, before a difficulty adjustment, the adversary’s mining strategy does not affect the average block generation time, and thus, $\mathbf{t}^{B}=1$. Note that while the canonical block generation time may differ, the average duration of each time step in our environment before a difficulty adjustment is determined by the average of both canonical and non-canonical block generation times, which remains unchanged. Therefore, the objective function before the difficulty adjustment can be defined as follows:
@@ F.2.1. Before the difficulty adjustment
F.2.2. After the difficulty adjustment
@@
where $\rho$ represents the time-averaged profit the adversary earns by following the optimal strategy, i.e., $\rho=O^{\mathcal{A}}(\pi^{*})$, where $\pi^{*}$ denotes the optimal strategy. At the start of training, the actual value of $\rho$ is unknown. Therefore, the training agents initialize with $\rho=O^{\mathcal{A}}(\pi^{\mathcal{H}})$, where $\pi^{\mathcal{H}}$ denotes the honest strategy. Certain test agents are tasked with measuring the time-averaged profit under the policy trained by the training agents at specific time steps during the training process. If this profit exceeds the highest time-averaged profit achieved so far, they prompt the training agents to update $\rho$ to the new highest time-averaged profit. For the detailed analysis of the non-linear objective function conversion readers are referred to (Sapirshtein et al., 2017). We briefly provide the intuition behind setting the time step reward as defined in equation 23 in the following. The value function of the initializing state $s_{0}$ under the time step reward introduced in equation 23 can be obtained as follows:
@@
F.3. Undercutting as a Continuous Action
In an A3C implementation, the neural network outputs a single scalar value representing the value of the current state, along with a probability distribution over the possible actions. For discrete actions, this probability distribution is a vector where each element represents the probability of selecting a specific action given the current state. For continuous actions, the network outputs the parameters of a probability distribution over the continuous action (Jia and Zhou, 2022). In our implementation, the neural network outputs a vector representing the probability distribution of the actions introduced in Section F.1. Additionally, it outputs two scalars representing the mean and variance of the duration of the undercutting action, which is modeled as a continuous action.
F.4. Implementation of the Undercutting Attack
@@
F.5. Implementation Details
Appendix G Weight-Time Equation
The estimated weight-time curves based on mempool statistics from December 17, 2023, are shown in Fig. 14. To depict Fig. 14, the transactions in mempool are categorized into five ranges of sat/vByte, with the base fee equal to $200$ sat/vByte. This base fee is chosen because, at all times on December 17, 2023, there was at least 1 virtual MB of transactions in the mempool paying a fee higher than $200$ sat/vByte. This indicates that, regardless of the block generation time, a block mined on this day could collect at least 200 million satoshis (2 BTC). An interesting observation from Fig. 14 is that, although the weights of transactions in all sat/vByte ranges are increasing over time, the rate of increase for weights in the lower sat/vByte ranges decreases over time (indicating a concave weight-time equation). In contrast, the rate of increase for weights in the higher sat/vByte ranges accelerates (indicating a convex weight-time equation). This observation aligns with the competition that exists in practice among transactions for inclusion in blocks. As the generation time for the next potential block increases, this competition intensifies, where newer transactions need to submit higher fees to maintain their chances of being included in the next block.
Appendix H Selfish Mining After Difficulty Adjustment
In Figure 15, the percentage increase in time-averaged profit of selfish mining after a difficulty adjustment is depicted as a function of the adversarial communication capability. This figure is depicted based on the following assumptions: i) the protocol reward is set to 00 BTC, ii) the mempool pattern observed on December 18, 2023, between 23:00 and 24:00 CET, has occurred, iii) the adversary’s mining share is equal to $\frac{1}{3}$.
@@ Appendix H Selfish Mining After Difficulty Adjustment
[…]
We divide canonical blocks mined under difficulty-adjusted selfish mining into groups $\{G_{1},G_{2},G_{3},\cdots\}$. A canonical block $B$ belongs to group $G_{i}$ for $i\in\{1,2,3,\cdots\}$ if the number of orphaned blocks mined between the parent of block $B$ and block $B$ itself is $i-1$. The average canonical block generation rate of blocks belonging to group $G_{i}$ is equal to $\frac{\lambda^{\prime}}{i}$. Let $a_{i}$ denote the ratio of the number of blocks belonging to group $G_{i}$ to the total number of canonical blocks, where $\sum_{i=1}^{\infty}{a_{i}}=1$. As the average canonical block generation rate under difficulty-adjusted selfish mining is also equal to $\lambda$, we have:
[…]
@@
[…]
@@
[…]
@@
[…]
@@
According to Theorem H.1, although the average canonical block generation rate is the same under both honest mining and difficulty-adjusted selfish mining, the total transaction fees distributed among canonical blocks under difficulty-adjusted selfish mining decrease due to the concavity of the time-fee function. This results in a lower profit percentage increase in the volatile model compared to honest mining.

### v3

@@ C.3. Strategy $\pi_{2}^{\text{Non-predictable}}$
Strategy $\pi_{2}^{\text{Non-predictable}}$ is presented in Table 3. Let $P_{0,0}$, $P_{0,1}$, $P_{i,i}$, $P_{i,i+1}$ denote the stationary probability of being in states $S_{0,0}$, $S_{0,1}$, $S_{i,i}$, $S_{i,i+1}$, respectively. We have:
(15) $\begin{split}&P_{0,0}=\frac{1-(1-\alpha)(\alpha+p)}{1-(1-p)\alpha(1-\alpha)},\quad P_{0,1}=\frac{(1-\alpha)p(1-\alpha(1-\alpha))}{1-(1-p)\alpha(1-\alpha)},\quad\\ &\sum_{i=1}^{\infty}{P_{i,i}}=P_{0,1}\frac{\alpha}{1-\alpha(1-\alpha)},\quad\sum_{i=1}^{\infty}{P_{i,i+1}}=P_{0,1}\frac{\alpha(1-\alpha)}{1-\alpha(1-\alpha)}.\end{split}$
Table 3. Strategy $\pi_{2}^{\text{Non-predictable}}$.
Initial State Action Probability Final State Reward
$S_{0,0}$
wait.
If the next block is adversarial, then switch to publish.
If the next block is honest with a normal transaction, then adopt.
$\alpha p$ $S_{0,0}$ $1+F$
$\alpha(1-p)$ $S_{0,0}$ 11
$(1-\alpha)p$ $S_{0,1}$ 0
$(1-\alpha)(1-p)$ $S_{0,0}$ 0
$S_{0,1}$
wait.
If the next block is honest with a normal transaction, then adopt
the whole honest fork.
If the next block is honest with a whale transaction, then adopt
except the last honest block.
$\alpha$ $S_{1,1}$ 0
$(1-\alpha)p$ $S_{0,1}$ 0
$(1-\alpha)(1-p)$ $S_{0,0}$ 0
$S_{i,i}$ $(i\geq 1)$
wait.
If the next block is adversarial, then publish.
$\alpha$ $S_{0,0}$ $1+F+i(1+pF)$
$(1-\alpha)$ $S_{i,i+1}$ 0
$S_{i,i+1}$ $(i\geq 1)$
wait.
If the next block is honest with a normal transaction, then adopt
the whole honest fork.
If the next block is honest with a whale transaction, then adopt
except the last honest block.
$\alpha$ $S_{i+1,i+1}$ 0
$(1-\alpha)p$ $S_{0,1}$ 0
$(1-\alpha)(1-p)$ $S_{0,0}$ 0
@@
The normalized numbers of canonical honest blocks, canonical adversarial blocks, and the normalized adversarial block reward achieved under strategy $\pi_{2}^{\text{Non-predictable}}$ can be obtained as follows:
@@
(16) $\begin{split}N_{\mathcal{H}}&=P_{0,0}(1-\alpha)(1-p)+P_{0,1}(1-\alpha)p+2P_{0,1}(1-\alpha)(1-p)\\ &\quad+\frac{P_{0,1}\alpha(1-\alpha)^{2}}{(1-\alpha(1-\alpha))^{2}}+\frac{P_{0,1}(2-p)\alpha(1-\alpha)^{2}}{1-\alpha(1-\alpha)}\enspace,\\ N_{\mathcal{A}}&=\left(\frac{P_{0,1}\alpha^{2}}{1-\alpha(1-\alpha)}\right)\left(\frac{1}{1-\alpha(1-\alpha)}+1\right)+P_{0,0}\alpha\enspace,\\ R_{\mathcal{A}}&=\left(\frac{P_{0,1}\alpha^{2}}{1-\alpha(1-\alpha)}\right)\left(\frac{(1+pF)}{1-\alpha(1-\alpha)}+(1+F)\right)+P_{0,0}\alpha(1+pF)\enspace.\end{split}$
Appendix D Simplified Mempool Environment: Supplementary
Deriving a time-fee equation compatible with the Bitcoin environment is challenging, as transaction fees vary across different periods due to various reasons such as fluctuations in Bitcoin’s supply and demand dynamics. This variability indicates that the optimal mining strategy for an adversary can differ significantly across different periods. To maximize its profit, the adversary must accurately predict the time-fee equation for the targeted attack period.
For the simplified mempool implementation, we extracted historical data (Blo, 2024) on total transaction fees and block generation times for Bitcoin blocks mined across different periods. Using regression techniques, we derived the time-fee equation for the target period. This time-fee equation enables us to estimate the reward function and analyze the profitability of mining strategies. As an example, in Figure 13, the total transaction fee of a block in BTC, denoted by fee, is depicted as a function of its generation time in minutes, denoted by $t$, for two different periods. Each point represents a block, and the line represents the time-fee equation. Figure 13(a) corresponds to Bitcoin blocks mined on January 1, 2023, with relatively low transaction fees per block, while Figure 13(b) corresponds to blocks mined on December 17, 2023, during a period of high transaction fees.
(a) January 1, 2023
(b) December 17, 2023
Figure 13. Linear time-fee equation.
As already mentioned in 4.2, a linear time-fee equation can be expressed as $\texttt{fee}(t)=\texttt{fee}_{0}+r_{\texttt{fee}}t$, where $\texttt{fee}_{0}$ and $r_{\texttt{fee}}$ are base fee and fee increase rate, respectively. At $t=0$, the transaction reward of the next potential block is equal to the base fee. This reflects reality, as there are always some low-value transactions in the mempool that miners can use to fill their blocks. As block generation time increases, the collected transaction fee rises by the fee increase rate for each additional unit of time due to the arrival of new transactions.
We assume that the time-fee equation is a linear function, and time is divided into $M\geq 2$ discrete steps of the form $\{t_{i}=i\Delta\mid i\in[M]\}$, $\Delta$ denotes the length of each time step. For a block with block generation time $t$, the transaction fee reward can be obtained as follows:
@@ Appendix E Limitations of the Simplified Environment
Non-linearity of the time-fee equation: In Figure 13 presented in Appendix D, a linear regression model is used to represent the time-fee equation. However, the block transaction fee does not necessarily increase linearly with the block generation time. This occurs when the rate of fee increase is not constant and decreases as the block generation time increases. In such cases, using alternative regression techniques rather than a linear model can provide a time-fee equation that better fits the behavior observed in the Bitcoin mempool. In Figure 14(a), we compare linear regression with a curve regression model given by $f(t)=0.6414\;\texttt{ln}(t+6.5209)-0.8419$ to derive the time-fee equation for Bitcoin blocks mined on May 3, 2023. The coefficient of determination (R-squared) for the curve regression is $0.6243$, while for the linear regression it is $0.5180$. This indicates that the curve regression provides a better fit compared to the linear regression, suggesting that the fee-increase rate diminishes with longer block generation times.
Dependence of fee rewards on the generation times of preceding blocks: Another technique that can lead to a more realistic mempool implementation is using multivariate regression instead of univariate regression. In univariate regression, the block transaction fee is estimated based solely on a single predictor, namely the block generation time. In contrast, multivariate regression allows us to account for additional predictors, such as the generation time of the parent block, in determining the block transaction fee. In Figure 14(b), we depict the regression plane for the block transaction fee on March 18, 2024, based on two predictors: the block generation time and the parent block generation time. The resulting time-fee equation is $f(t_{1},t_{2})=0.1697+0.0079*t_{1}+0.0046*t_{2}$, where $t_{1}$ and $t_{2}$ represent the block generation time and the parent block generation time, respectively. Note that if the parent block generation time is longer, more fee-valuable transactions will likely remain in the mempool after the parent block is mined, allowing the next block to collect these transactions. As the time-fee equation shows, the fee increase rate with respect to the block generation time is higher than the fee increase rate with respect to the parent block generation time. This implies that the block generation time is a more effective predictor of the transaction fee compared to the parent block generation time.
@@
Figure 14. Time-fee equation.
@@ Appendix F A3C Implementation
F.1. Objective Function
@@
F.1.1. Before the difficulty adjustment
As discussed in Section 2.2.1, before a difficulty adjustment, the adversary’s mining strategy does not affect the average block generation time, and thus, $\mathbf{t}^{B}=1$. Note that while the canonical block generation time may differ, the average duration of each time step in our environment before a difficulty adjustment is determined by the average of both canonical and non-canonical block generation times, which remains unchanged. Therefore, the objective function before the difficulty adjustment can be defined as follows:
@@ F.1.1. Before the difficulty adjustment
F.1.2. After the difficulty adjustment
@@
where $\rho$ represents the time-averaged profit the adversary earns by following the optimal strategy, i.e., $\rho=O^{\mathcal{A}}(\pi^{*})$, where $\pi^{*}$ denotes the optimal strategy. At the start of training, the actual value of $\rho$ is unknown. Therefore, the training agents initialize with $\rho=O^{\mathcal{A}}(\pi^{\mathcal{H}})$, where $\pi^{\mathcal{H}}$ denotes the honest strategy. Certain test agents are tasked with measuring the time-averaged profit under the policy trained by the training agents at specific time steps during the training process. If this profit exceeds the highest time-averaged profit achieved so far, they prompt the training agents to update $\rho$ to the new highest time-averaged profit. For the detailed analysis of the non-linear objective function conversion, readers are referred to (Sapirshtein et al., 2017). We briefly provide the intuition behind setting the time step reward as defined in equation 23 in the following. The value function of the initializing state $s_{0}$ under the time step reward introduced in equation 23 can be obtained as follows:
@@
F.2. Implementation Details
F.3. Implementation of the Undercutting Attack
@@
Appendix G Selfish Mining After Difficulty Adjustment
In Figure 15, the percentage increase in time-averaged profit of selfish mining after a difficulty adjustment is depicted as a function of the adversarial communication capability. This figure is depicted based on the following assumptions: i) the protocol reward is set to 0 BTC, ii) the mempool pattern observed on December 18, 2023, between 23:00 and 24:00 CET, has occurred, iii) the adversary’s mining share is equal to $\frac{1}{3}$.
@@ Appendix G Selfish Mining After Difficulty Adjustment
(26) $r_{\mathcal{H}}=\int_{0}^{\infty}f(t)\lambda e^{-\lambda t}\,dt\enspace.$
We divide canonical blocks mined under difficulty-adjusted selfish mining into groups $\{G_{1},G_{2},G_{3},\cdots\}$. A canonical block $B$ belongs to group $G_{i}$ for $i\in\{1,2,3,\cdots\}$ if the number of orphaned blocks mined between the parent of block $B$ and block $B$ itself is i1i-1. The average canonical block generation rate of blocks belonging to group $G_{i}$ is equal to $\frac{\lambda^{\prime}}{i}$. Let $a_{i}$ denote the ratio of the number of blocks belonging to group $G_{i}$ to the total number of canonical blocks, where $\sum_{i=1}^{\infty}{a_{i}}=1$. As the average canonical block generation rate under difficulty-adjusted selfish mining is also equal to $\lambda$, we have:
(27) $\frac{1}{\lambda}=\sum_{i=1}^{\infty}{a_{i}\int_{0}^{\infty}t^{\prime}\frac{\lambda^{\prime}}{i}e^{-\frac{\lambda^{\prime}}{i}t^{\prime}}\,dt^{\prime}}\enspace.$
@@
(29) $r_{\mathcal{S}}=\sum_{i=1}^{\infty}{a_{i}\int_{0}^{\infty}f(t^{\prime})\frac{\lambda^{\prime}}{i}e^{-\frac{\lambda^{\prime}}{i}t^{\prime}}\,dt^{\prime}}\enspace.$
@@
(30) $r_{\mathcal{S}}=\sum_{i=1}^{\infty}{a_{i}\int_{0}^{\infty}f(\frac{\lambda}{\lambda^{\prime}}it)\lambda e^{-\lambda t}\,dt}\enspace.$
@@
(31) $r_{\mathcal{S}}<\int_{0}^{\infty}f(\frac{\lambda}{\lambda^{\prime}}\big{(}\sum_{i=1}^{\infty}{a_{i}i}\big{)}t)\lambda e^{-\lambda t}\,dt\enspace.$
@@
According to Theorem G.1, although the average canonical block generation rate is the same under both honest mining and difficulty-adjusted selfish mining, the total transaction fees distributed among canonical blocks under difficulty-adjusted selfish mining decrease due to the concavity of the time-fee function. This results in a lower profit percentage increase in the volatile model compared to honest mining.
Appendix H Selfish Mining Before Difficulty Adjustment: Supplementary
Figure 9 shows the percentage increase in the adversary’s time-averaged profit before a difficulty adjustment resulting from conducting selfish mining. This figure is depicted based on different ratios of petty-compliant miners, over the mempool patterns observed on two different days: January 1, 2024, and March 1, 2024, both between 8:00 and 9:00 AM CET. The figure is generated using the following configuration parameters: i) the protocol reward is set to 0 BTC, and ii) the adversary’s mining share and communication capability are set to $\frac{1}{3}$ and $50\%$, respectively.
Figure 17. The percentage increase in time-averaged profit.
The greater percentage increase observed under the mempool pattern of March 1, 2024, can be attributed to a higher normalized rate of growth in fee rewards during this period. Specifically, the ratio of the average fee reward for a block generated in 20 minutes to that for a block generated in 10 minutes is 1.29 under the mempool pattern of January 1, 2024, and 1.4 under that of March 1, 2024. This indicates that if the adversary succeeds in orphaning a preceding block and thereby extends the block generation time for its own block, it can accumulate a relatively higher fee reward under the mempool pattern of March 1, 2024.